A network hardware failure took the BHRIC’s Virtual Desktops offline on Saturday the 5th. We have resolved the issue and are currently testing services to make sure they are online.
BHRIC Virtual Desktop outage
Account Self Service Portal Emergency Maintenance
An emergency maintenance is being performed to correct an issue related to Account Self Service and password resets.
During this time you may see a delay in password resets. If you see any errors please retry at a later time. If you continue to have issues please contact the Help Desk for further assistance (561) 297-3999
We apologize for any inconvenience this may cause.
[Resolved] Research Computing Data Transfer nodes offline
The bellow issue has been resolved and the service restored.
[RESOLVED] Due to a software issue on our data transfer nodes this service is temporarily offline. We are rebuilding the Windows file share service and hope to have it online by the end of the day.
[Resolved] BHRIC Virtual Desktop Infrastructure downtime
The Biomedical Health Research Informatics Core experienced an error that prevented virtual desktops being rebooted, started, or migrated. The error occurred between 7:00 – 7:50 am August 22 and appears to have been resolved. For more information please contact the Research Computing team.
Owl Apps login issue: Google Services “Something went wrong”
We are investigating a potential issue with signing into Google services.
We will provide more information by Monday, 2019-08-19 12:45 US/Pacific.
How to diagnose:
Users may have issues signing into Google services. They may receive the error “Something went wrong. Sorry, something went wrong there. Try again.”
There’s no workaround available at this time, but we’ll let you know if we learn of one.
[RESOLVED] Starfish – Down
Starfish is currently down and we are working with Hobsons to resolve the issue.
BHRIC Health System Experiencing Issues
Following this weekend’s network updates, the BHRIC Health System is experiencing connectivity issues. Users may have difficulty logging in to VDIs and Servers. The team is working diligently to restore services to normal operation. We will update as progress continues.